# Efficient and operational quantifier of divisibility in terms of channel discrimination

### v1

@@ Abstract
The understanding of open quantum systems is crucial for the development of quantum technologies. Of particular relevance is the characterisation of divisible quantum dynamics, seen as a generalisation of Markovian processes to the quantum setting. Here, we propose a way to detect divisibility and quantify how non-divisible a quantum channel is through the concept of channel discrimination. We ask how well we can distinguish generic dynamics from divisible dynamics. We show that this question can be answered efficiently through semidefinite programming, which provides us with an operational and efficient way to quantify divisibility.
@@ I Introduction
No physical system is completely isolated from its surrounding environment. The unavoidable interaction between a system and its environment renders the possibility of scaling quantum phenomena to the macroscopic world Zurek (2003); Blume-Kohout and Zurek (2006); Schlosshauer (2005) and represents the main barrier for practical developments of new quantum technologies Lidar et al. (1998); Arndt and Hornberger (2014). Unless one counteracts the detrimental effects of decoherence, which typically accumulate exponentially fast with the system’s size Aolita et al. (2008), the advantages provided by quantum information processing with respect to classical strategies become unfeasible in practice. Because of this, a great effort is devoted to understanding and modeling decoherence processes and proposing methods that counteract their effect, such as noise mitigation Wallman and Emerson (2016); Berg et al. (2022); Cai et al. (2023) and quantum error correction Lidar and Brun (2013) schemes.
@@
Formally, the system-environment interaction causes the system’s evolution to be non-unitary, i.e., to be described by a more general, completely positive, and trace-preserving (CPTP) map. Given quantum dynamics, a natural question is whether it is Markovian or not Ángel Rivas et al. (2014); Breuer et al. (2016). Unlike the classical case, where Markovian concepts of memory effects and divisibility are linked, different notions of Markovianity arise in quantum mechanics Chruściński et al. (2011); Ángel Rivas et al. (2014). This has opened up an active field for discussions on how to properly define quantum Markovianity. Several proposals have defined Markovian processes, including channel divisibility Rivas et al. (2010), non-increase of state distinguishability Breuer et al. (2009); Buscemi and Datta (2016); Bae and Chruściński (2016), non-increase of system-environment correlations De Santis et al. (2019, 2020); Kołodyński et al. (2020); Abiuso et al. (2022), among others Pollock et al. (2018); Budini (2018); Capela et al. (2020, 2022).
@@
In this paper, we focus on the notion of divisibility and ask the following questions: (i) Can we efficiently detect whether a family of quantum channels is divisible? (ii) Given a non-divisible quantum dynamics, can we quantify its degree of non-divisibility in an efficient and meaningful operational way? We provide a positive answer to both questions by proposing quantifying divisibility as the minimum distance from the channel to a divisible one. We use a distance induced by the diamond norm between channels as a measure of distinguishability. Such distance can be calculated via semidefinite programming so that efficient numerical methods can be used for its calculation Skrzypczyk and Cavalcanti (2023). Moreover, the distance induced by the diamond norm also has an operational interpretation regarding the minimum error that one can make in a channel discrimination protocol when we use it for general probe states and measurements Pirandola et al. (2019); Brandão et al. (2015). Using this quantifier, we investigate the amount of non-divisibility present in two different models. The first one is a versatile toy model for studying open quantum system dynamics: the collisional model Ciccarello et al. (2022). The second one is a paradigmatic decoherence channel: the dephasing model.
@@
Note that in the context of a resource theory of Markovianity, channel discrimination and semidefinite programming for robustness have already been explored in Ref. Anand and Brun (2019a). However, unlike the analysis presented here, the framework investigated in Ref. Anand and Brun (2019a) is only immediately applicable when the underlying set of Markovian dynamics forms a convex set. Moreover, the scheme proposed here is valid for all non-Markovian dynamics (invertible or not), which sets our work apart from universal non-Markovian quantifiers based on correlations De Santis et al. (2019, 2020); Kołodyński et al. (2020).
@@ II Divisibility
The scenario we study in the paper is schematically depicted in Fig. 1. We investigate quantum dynamics that evolve quantum systems from time $t_{0}$ to time $t_{N}$. There is a paradigmatic way in which we can analyze these dynamics: by cutting the time interval $[t_{0},t_{N}]$ into $N$ steps and reconstructing (by means of quantum process tomography Mohseni et al. (2008)) the maps that describe the time evolution from $t_{0}$ to each $t_{k}$ $(k=1,\cdots,N)$. Mathematically, each of these maps is described by a CPTP map
@@
where $\mathcal{L}(\mathcal{H})$ represents the set of linear operators acting on the Hilbert space $\mathcal{H}$. In other words, we are given a stroboscopic characterization of the dynamics in terms of the family of maps :=$\mathcal{F}:=\{\mathcal{N}_{k|0}\}_{k=1}^{N}$, each of which representing the time evolution of a quantum system from an initial time step $t_{0}$ to a given time step $t_{k}$.
@@
In this case, we simply need to check whether $\mathcal{N}_{2|1}$ is CP, P, or neither. Although there are cases where this calculation does not bring any complication, it does not solve two problems: (i) what if $\mathcal{N}_{1|0}$ is not invertible? (ii) Given that the dynamics is not divisible, how can we quantify its degree of non-divisibility? (iii) Lastly, what if $\mathcal{N}_{1|0}$ is invertible, but it is the map representing a large many-body quantum system’s dynamics, and, accordingly, determining its inverse is not an easy task either? In the next sections, we present an SDP formulation that is able to solve these questions.
@@ III Conditional Quantum States and Divisibility - A Brief Review
Here, we review the two mathematical tools we will use extensively in this work. Inspired by the approach developed in Refs. Leifer and Spekkens (2013, 2014) we start this section with a brief review of the Choi-Jamiołkowski (CJ) isomorphism. Next, we carefully define what we mean by divisibility for a given quantum dynamics—in particular, the completely positive and trace-preserving CP-divisibility and P-divisibility. We conclude this section by connecting divisibility with CJ-states, paving the road for the task of designing an efficient witness for (CP or P) divisibility. For completeness, we provide the proofs of the propositions in Appendix A.
@@ III.1 Conditional States Approach
We start defining the Jamiołkoswki isomorphism Jamioł kowski (1972). In our work, it assumes the following form.
@@
$\mathcal{N}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{B})$ (5)
@@
$\rho:=(\mbox{id}_{A}\otimes\mathcal{N}\circ\mbox{T}_{A})\sum_{i,j}^{d}\ket{i}\bra{j}_{A}\otimes\ket{i}\bra{j}_{A},$ (6)
where $d=\mbox{dim}(\mathcal{H}_{A})$ and the transposition $\mbox{T}_{A}$ is taken with respect to some basis in $\mathcal{H}_{A}$. On the other hand, the action of $\mathcal{N}$ on $\mathcal{L}(\mathcal{H}_{A})$ is given by
$\mathcal{N}(\sigma_{A})=\mbox{Tr}_{A}[\rho(\sigma_{A}\otimes\mathds{1}_{B})],$ (7)
where $\mathcal{N}(\sigma_{A})\in\mathcal{D}(\mathcal{H}_{B})$ and the set $\mathcal{D}(\mathcal{H}_{X})=\{\rho\in\mathcal{L}(\mathcal{H}_{X})|\rho\geq 0,\text{Tr}(\rho)=1\}$ is the state space over $\mathcal{H}_{X}$.
We emphasise that contemporary literature is still debating about whether the transposition map should or should not appear in Eq. (6)—see Refs. Milz et al. (2019); Oreshkov et al. (2012); Chiribella et al. (2009). Different authors are more inclined towards one or another, but in this work, we will stick to the def. 7 above, following Refs. Leifer and Spekkens (2013, 2014).
The fact that $\rho$ and $\mathcal{N}$ are isomorphically connected can be easily checked, since
$\mbox{Tr}_{A}[\rho(\sigma_{A}\otimes\mathds{1}_{B})]=$
$=\mbox{Tr}_{A}\left[(\mbox{id}\otimes\mathcal{N})(\sum_{i,j}^{d}\ket{i}\bra{j}\otimes\ket{j}\bra{i})(\sigma_{A}\otimes\mathds{1}_{B})\right]$ (8)
@@
$=\sum_{i,j}^{d}(\sigma_{A})_{j,i}\mathcal{N}(\ket{j}\bra{i})=\mathcal{N}(\sigma_{A}).$
@@
Let $\mathcal{N}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{B})$ be a linear map, and let $\rho\in\mathcal{D}(\mathcal{H}_{B}\otimes\mathcal{H}_{A})$ be the Jamiołkowski isomorphic operator associated to it. It follows that $\rho$ satisfies
@@
    $\mbox{Tr}_{B}[\rho]=\mathds{1}_{A}$
if, and only if, $\mathcal{N}\circ\mbox{T}_{A}$ is a completely positive and trace preserving map.
Our starting point will generally be quantum channels rather than arbitrary linear maps. We will be interested in determining precisely when a given collection $\{\Lambda_{n}\}_{n\in[N]}$ of CPTP maps are divisible. In this case, Prop. III.2 remains applicable, as it suffices to change item (a) from demanding that $\rho^{T_{A}}\geq 0$ to demanding, instead, that $\rho\geq 0$. So whenever working with CPTP maps, to factor in the action of the partial transposition, instead of dealing with $\rho$, we will consider a slightly different operator. We will use $\varrho:=\rho^{T_{A}}\geq 0$, as it results in a positive operator. In summary, $\rho$ is the bipartite conditional state associated with $\mathcal{N}$, and $\varrho$ is the bipartite conditional state that must be positive to render $\mathcal{N}$ completely positive.
Throughout this work, we will consistently use a particular index notation for states and channels alike. We will write time steps and Hilbert spaces’ labels in a manner reminiscent of the “given” notation commonly used for conditional probabilities. Thus, when a channel is written as $\mathcal{N}_{j|i}$, then it is meant to imply, first, that
$\mathcal{N}_{j|i}:\mathcal{L}(\mathcal{H}_{i})\rightarrow\mathcal{L}(\mathcal{H}_{j}),$ (9)
and, secondly, that thought as an evolution map, it determines the dynamics from time step $t_{i}$ to time step $t_{j}$—with $t_{j}\geq t_{i}$. Analogously, the Choi-Jamiołkowski state associated with this channel will be written as $\varrho_{j|i}$, where
$\varrho_{j|i}\in\mathcal{L}(\mathcal{H}_{i})\otimes\mathcal{L}(\mathcal{H}_{j}).$ (10)
This notation improves the calculations and the reading of our expressions. This doubled notation also facilitates the comparison with other works where expressions like $j|i$ have a Bayesian-probabilistic meaning—see Refs. Leifer and Spekkens (2013, 2014).
Definition 7 shows how to connect a single quantum channel $\mathcal{N}_{i|j}$ with its respective conditional state $\varrho_{i|j}$. It is quite natural to ask how we can calculate the composition of two channels. In other words, given $\mathcal{N}_{i|j}$ and $\mathcal{N}_{k|i}$ we may want to determine what the Jamiołkowski image of the composition $\mathcal{N}_{k|i}\circ\mathcal{N}_{i|j}$ is. The proposition below, whose proof is in Appendix B, address this question.
@@
Let $\varrho_{2|1}$, $\varrho_{3|1}$, $\varrho_{3|2}$ be the Choi-Jamiołkowski images of the CPTP maps $\mathcal{N}_{2|1}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{B})$, $\mathcal{N}_{3|1}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{C})$ and $\mathcal{N}_{3|2}:\mathcal{L}(\mathcal{H}_{B})\rightarrow\mathcal{L}(\mathcal{H}_{C})$ respectively. The composition
$\mathcal{N}_{3|1}=\mathcal{N}_{3|2}\circ\mathcal{N}_{2|1}$ (11)
@@
$\varrho_{3|1}=\mbox{Tr}_{B}[(\mathds{1}_{A}\otimes\varrho_{3|2})^{T_{B}}(\varrho_{2|1}\otimes\mathds{1}_{C})].$ (12)
Whenever we have to decide whether a dynamics =$\mathcal{F}=\{\mathcal{N}_{k|0}\}_{k\in[N]}$ is divisible (either P or CP), Prop. 12 tell us it suffices to check out whether an equation like Eq. (12) holds true for each pair of time steps $(t_{i},t_{i+1})$. By doing so, instead of looking for the existence of intermediate CPTP or PTP maps verifying the composition in Eq. (11), what can be rather difficult to establish  Buscemi and Datta (2016); Ángel Rivas et al. (2014); Filippov et al. (2017), we will show that problem can be addressed by checking $N-1$ instances of an easier optimisation problem.
To establish our main result—an efficient quantifier of divisibility in terms of channel discrimination—we will, for simplicity, focus on only two intermediate time steps: $t_{i}$ and $t_{i+1}$. The following corollary is a direct rewriting of Prop. 12, more adequate to our framework.
@@
$\varrho_{i+1|0}=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\varrho_{i+1|i})^{T_{B}}(\varrho_{i|0}\otimes\mathds{1}_{C})\right],$ (14)

### v2

# Efficient and operational quantifier of non-divisibility in terms of channel discrimination

@@ Abstract
The understanding of open quantum systems is crucial for the development of quantum technologies. Of particular relevance is the characterisation of divisible quantum dynamics, seen as a generalisation of Markovian processes to the quantum setting. Here, we propose a way to detect divisibility and quantify how non-divisible a quantum channel is through the concept of channel discrimination. We ask how well we can distinguish generic dynamics from divisible dynamics. We show that this question can be answered efficiently through semidefinite programming, which provides us with an operational and efficient way to quantify non-divisibility.
@@ I Introduction
No physical system is completely isolated from its surrounding environment. The unavoidable interaction between a system and its environment renders the possibility of scaling quantum phenomena to the macroscopic world [1, 2, 3] and represents the main barrier for practical developments of new quantum technologies [4, 5]. Unless one counteracts the detrimental effects of decoherence, which typically accumulate exponentially fast with the system’s size [6], the advantages provided by quantum information processing with respect to classical strategies become unfeasible in practice. Because of this, a great effort is devoted to understanding and modeling decoherence processes and proposing methods that counteract their effect, such as noise mitigation [7, 8, 9] and quantum error correction [10] schemes.
@@
Formally, the system-environment interaction causes the system’s evolution to be non-unitary, i.e., to be described by a more general, completely positive, and trace-preserving (CPTP) map. Given quantum dynamics, a natural question is whether it is Markovian or not [11, 12]. Unlike the classical case, where Markovian concepts of memory effects and divisibility are linked, different notions of Markovianity arise in quantum mechanics [13, 11]. This has opened up an active field for discussions on how to properly define quantum Markovianity. Several proposals have defined Markovian processes, including channel divisibility [14], non-increase of state distinguishability [15, 16, 17], non-increase of system-environment correlations [18, 19, 20, 21, 22], among others [23, 24, 25, 26, 27, 28].
@@
In this paper, we focus on the notion of divisibility and ask the following questions: (i) Can we efficiently detect whether a family of quantum channels is divisible? (ii) Given a non-divisible quantum dynamics, can we quantify its degree of non-divisibility in an efficient and meaningful operational way? We provide a positive answer to both questions by proposing quantifying non-divisibility as the minimum distance from the channel to a divisible one. We use a distance induced by the diamond norm between channels as a measure of distinguishability. Such distance can be calculated via semidefinite programming so that efficient numerical methods (i.e. polynomial in the dimension of the problem) can be used for its calculation [29]. Moreover, the distance induced by the diamond norm also has an operational interpretation regarding the minimum error one can make in a channel discrimination protocol when we use it for general probe states and measurements [30, 31]. We use this quantifier to investigate the amount of non-divisibility present in three different models. The first is a versatile toy model for studying open quantum system dynamics: the collisional model [32]. The second is a paradigmatic decoherence channel: the dephasing model. Finally, the third model explores a convex combination of unitary channels.
@@
Note that in the context of a resource theory of Markovianity, channel discrimination and semidefinite programming for robustness have already been explored in Ref. [33]. However, unlike the analysis presented here, the framework investigated in Ref. [33] is only immediately applicable when the underlying set of Markovian dynamics forms a convex set. Moreover, the scheme proposed here is valid for all non-Markovian dynamics (invertible or not), which sets our work apart from universal non-Markovian quantifiers based on correlations [19, 20, 21]. Similarly, recently we became aware of yet another novel measure of non-Markovianity, one that explores the relationship between CP indivisibility and the compatibility of quantum channels [34, 35].
@@ II Divisibility
The scenario we study in the paper is schematically depicted in Fig. 1. We investigate quantum dynamics that evolve quantum systems from time $t_{0}$ to time $t_{N}$. There is a paradigmatic way in which we can analyze these dynamics: by cutting the time interval $t_{0},t_{N}]$ into $N$ steps and reconstructing (by means of quantum process tomography [36]) the maps that describe the time evolution from $t_{0}$ to each $t_{k}$ $(k=1,\cdots,N)$. Mathematically, each of these maps is described by a CPTP map
@@
where each $\mathcal{L}(\mathcal{H}_{k})$ represents the set of linear operators acting on the Hilbert space $\mathcal{H}_{k}$, and each $\mathcal{H}_{k}$ is associated with an instant of time. In other words, we are given a stroboscopic characterization of the dynamics in terms of the family of maps $\mathcal{F}:=\{\mathcal{N}_{k|0}\}_{k=1}^{N}$, each of which representing the time evolution of a quantum system from an initial time step $t_{0}$ to a given time step $t_{k}$. The fact that there is an apparent mismatch between the sub-index location ($k|0$) and the domain ($\mathcal{L}(\mathcal{H}_{0})$), contra-domain $(\mathcal{L}(\mathcal{H}_{k}))$ is explained in def. 7 below. For now, the object $\mathcal{N}_{k|0}$ should be read as the map governing the dynamics at $t_{k}$ given $t_{0}$.
@@
In this case, we simply need to check whether $\mathcal{N}_{2|1}$ is CP, P, or neither. Although there are cases where this calculation does not bring any complication, it does not solve two problems: (i) what if $\mathcal{N}_{1|0}$ is not invertible? (ii) Given that the dynamics is not divisible, how can we quantify its degree of non-divisibility? In the next sections, we present an SDP formulation that is able to solve these questions.
@@ III Conditional Quantum States and Divisibility - A Brief Review
Here, we review the two mathematical tools we will use extensively in this work. Inspired by the approach developed in Refs. [37, 38] we start this section with a brief review of the Choi-Jamiołkowski (CJ) isomorphism. Next, we carefully define what we mean by divisibility for a given quantum dynamics—in particular, the completely positive and trace-preserving CP-divisibility and P-divisibility. We conclude this section by connecting divisibility with CJ-states, paving the road for the task of designing an efficient witness for (CP or P) divisibility. For completeness, we provide the proofs of the propositions in Appendix A.
@@ III.1 Conditional States Approach
We start defining the Jamiołkoswki isomorphism [39]. In our work, it assumes the following form.
@@
$\mathcal{N}_{B|A}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{B})$ (5)
@@
where $d=\mbox{dim}(\mathcal{H}_{A})$ and the transposition $\mbox{T}_{A}$ is taken with respect to some basis in $\mathcal{H}_{A}$. On the other hand, the action of $\mathcal{N}_{B|A}$ on $\mathcal{L}(\mathcal{H}_{A})$ is given by
$\mathcal{N}_{B|A}(\sigma_{A})=\mbox{Tr}_{A}\left[\rho_{B|A}(\sigma_{A}\otimes\mathds{1}_{B})\right],$ (7)
where $\mathcal{N}_{B|A}(\sigma_{A})\in\mathcal{D}(\mathcal{H}_{B})$ and the set $\mathcal{D}(\mathcal{H}_{B}):=\{\rho_{B}\in\mathcal{L}(\mathcal{H}_{X})|\rho_{B}\geq 0,\text{Tr}(\rho_{B})=1\}$ is the state space over $\mathcal{H}_{B}$.
Remark: As we mentioned before, the notation $\mathcal{N}_{B|A}$ is intended to mean a map that outputs a state on a Hilbert space $\mathcal{H}_{B}$ given that its input Hilbert space is $\mathcal{H}_{A}$. We adhere to this ‘$B$ given $A$’ notation to create a closer parallel with refs. [37, 38]. We will show how to adapt this notation to our case (with time stamps) in a minute.
We emphasise that contemporary literature is still debating about whether the transposition map should or should not appear in Eq. (6)—see Refs. [40, 41, 42]. Different authors are more inclined towards one or another, but in this work, we will stick to the def. 7 above, following Refs. [37, 38]. In any case, the fact that $\rho_{B|A}$ and $\mathcal{N}_{B|A}$ are isomorphically connected can be easily checked, since
$\mbox{Tr}_{A}[\rho_{B|A}(\sigma_{A}\otimes\mathds{1}_{B})]=$
[…]
@@
Let $\mathcal{N}_{B|A}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{B})$ be a linear map, and let $\rho\in\mathcal{D}(\mathcal{H}_{B}\otimes\mathcal{H}_{A})$ be the Jamiołkowski isomorphic operator associated to it. It follows that $\rho$ satisfies
@@
    $\mbox{Tr}_{B}[\rho_{B|A}]=\mathds{1}_{A}$
if, and only if, $\mathcal{N}_{B|A}\circ\mbox{T}_{A}$ is a completely positive and trace preserving map.
Our starting point will generally be quantum channels rather than arbitrary linear maps. We will be interested in determining precisely when a given collection $\{\mathcal{N}_{t|0}\}_{t\in[N]}$ of CPTP maps are divisible. In this case, Prop. III.2 remains applicable, as it suffices to change item (a) from demanding that $\rho^{T_{A}}_{t|0}\geq 0$ to demanding, instead, that $\rho_{t|0}\geq 0$. So whenever working with CPTP maps, to factor in the action of the partial transposition, instead of dealing with $\rho_{t|0}$, we will consider a slightly different operator. We will use $\varrho_{t|0}:=\rho^{T_{A}}_{t|0}\geq 0$, as it results in a positive operator. In summary, $\rho_{t|0}$ is the bipartite conditional state associated with $\mathcal{N}_{t|0}$, and $\varrho_{t|0}$ is the bipartite conditional state that must be positive to render $\mathcal{N}_{t|0}$ completely positive.
Throughout this work, we will consistently use a particular index notation for states and channels alike. We will write time steps and Hilbert spaces’ labels in a manner reminiscent of the “given” notation commonly used for conditional probabilities. Thus, when a channel is written as $\mathcal{N}_{t_{j}|t_{i}}$, then it is meant to imply, first, that
$\mathcal{N}_{t_{j}|t_{i}}:\mathcal{L}(\mathcal{H}_{i})\rightarrow\mathcal{L}(\mathcal{H}_{j}),$ (9)
and, secondly, that thought as an evolution map, it determines the dynamics from time step $t_{i}$ to time step $t_{j}$—with $t_{j}\geq t_{i}$. Analogously, the Choi-Jamiołkowski state associated with this channel will be written as $\varrho_{t_{j}|t_{i}}$, where
$\varrho_{t_{j}|t_{i}}\in\mathcal{L}(\mathcal{H}_{i})\otimes\mathcal{L}(\mathcal{H}_{j}).$ (10)
This notation improves the calculations and the reading of our expressions. This doubled notation also facilitates the comparison with other works where expressions like $a|b$ have a Bayesian-probabilistic meaning—see Refs. [37, 38].
Definition 7 shows how to connect a single quantum channel $\mathcal{N}_{t_{i}|t_{j}}$ with its respective conditional state $\varrho_{t_{i}|t_{j}}$. It is quite natural to ask how we can calculate the composition of two channels. In other words, given $\mathcal{N}_{t_{i}|t_{j}}$ and $\mathcal{N}_{t_{k}|t_{i}}$ we may want to determine what the Jamiołkowski image of the composition $\mathcal{N}_{t_{k}|t_{i}}\circ\mathcal{N}_{t_{i}|t_{j}}$ is. The proposition below, whose proof is in Appendix B, address this question.
@@
Let $\varrho_{1|0}$, $\varrho_{2|0}$, $\varrho_{2|1}$ be the Choi-Jamiołkowski images of the CPTP maps $\mathcal{N}_{1|0}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{1})$, $\mathcal{N}_{3|1}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{2})$ and $\mathcal{N}_{2|1}:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{2})$ respectively. The composition
$\mathcal{N}_{2|0}=\mathcal{N}_{2|1}\circ\mathcal{N}_{1|0}$ (11)
@@
$\varrho_{2|0}=\mbox{Tr}_{\mathcal{H}_{1}}[(\mathds{1}_{\mathcal{H}_{0}}\otimes\varrho_{2|1})^{T_{\mathcal{H}_{1}}}(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})].$ (12)
Whenever we have to decide whether a dynamics $\mathcal{F}=\{\mathcal{N}_{k|0}\}_{k\in[N]}$ is divisible (either P or CP), Prop. 12 tell us it suffices to check out whether an equation like Eq. (12) holds true for each pair of time steps $(t_{i},t_{i+1})$. By doing so, instead of looking for the existence of intermediate CPTP or PTP maps verifying the composition in Eq. (11), what can be rather difficult to establish  [16, 11, 43], we will show that problem can be addressed by checking N1N-1 instances of an easier optimisation problem.
To establish our main result—an efficient quantifier of non-divisibility in terms of channel discrimination—we will, for simplicity, focus on only two intermediate time steps: $t_{i}$ and $t_{i+1}$. The following corollary is a direct rewriting of Prop. 12, more adequate to our framework.
@@
$\varrho_{i+1|0}=\mbox{Tr}_{\mathcal{H}_{i}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\varrho_{i+1|i})^{T_{\mathcal{H}_{i}}}(\varrho_{i|0}\otimes\mathds{1}_{\mathcal{H}_{i+1}})\right],$ (14)